OpenStudy (anonymous):

if tan(theta)=3/5 and sin(theta)<0 how do i find sin?

OpenStudy (anonymous):

find theta first. find this my taking the inverse tangent of 3/5

OpenStudy (anonymous):

take the inverse of tangent 3/5?

OpenStudy (anonymous):

\[\sin \left(-\tan ^{-1}\left(\frac{3}{5}\right)\right)=-\frac{3}{\sqrt{34}} \]

OpenStudy (anonymous):

ok sin is -3/sqrt(34)